# Cross-Entropy Games for Language Models: From Implicit Knowledge to General Capability Measures

### v1

@@ 6.1.1. Web Exploration Intuition
If we assume that space of Xent Games is $\alpha$-connected for some $\alpha>0$, we can think of it as an (infinite) graph, we can put a downstream link from $G_{1}$ to $G_{2}$ if $G_{1}\rightmoon_{\alpha}G_{2}$ and an upstream link if $G_{1}\leftmoon_{\alpha}G_{2}$. We can hence leverage this structure to think of the space of games as a kind of ‘inter-game web’ and use this vision as a guiding principle to explore it like a (very) complex network.
Note that unlike the internet, it is a priori very non-trivial to list all the links from a given game; however, we can think that in practice, there is a from a finite list of games $\Sigma$, it is definitely possible to determine which ones are $\alpha$-linked and which ones are not, thus equipping $\Sigma$ with that structure.
@@
Following this principle, it is tempting to think of the exploration of the space of games similar to ‘random web browsing’ (following the philosophy of the PageRank algorithm [PBRW99]). If in some sense one wants to find a ‘good sample’ of the internet, one can browse randomly and save the visited pages on the fly. Depending on the application we may even move for long enough so our visit becomes independent from the starting point (note also that the longer we ‘walk’ the more biased towards the more connected nodes we get). From this, we could in principle use various heuristics (e.g. the PageRank algorithm) to rank the games that we care about for a target application.
@@ 6.1.2. Toy Model
  • For any pair of games $G,\tilde{G}$, corresponding to $\left(g_{1},g_{2},\ldots,g_{n},\ldots\right)$ and $\left(\tilde{g}_{1},\tilde{g}_{2},\ldots,\tilde{g}_{n},\ldots\right)$, we define $\mathcal{V}_{G}\left(\tilde{G}\right)$ as $\frac{1}{\|\tilde{G}\|_{1}}\sum_{i=1}^{\infty}\min\left(\tilde{g}_{i},g_{i}\right)$, i.e. which fraction of $\tilde{g}$ skills are ‘captured’ by playing $g$ and $\mathcal{V}_{\tilde{G}}\left(G\right)$ as $\frac{1}{\|G\|_{1}}\sum_{i=1}^{\infty}\min\left(g_{i},\tilde{g}_{i}\right)$ the symmetric quantity.
@@
This simply suggests the (naive, but useful) view that games are associated with skills that are entries in a vector, and that learning these amounts to acquire skills proportionally to the vector of skills.
@@ 6.2.1. Meta-Game: Informal Evolution Dynamics Idea
In a sense, the question of measuring ‘general capabilities’ of models comes down to a motivation: why do we want to evaluate general abilities, in the first place? A plausible answer is because agents will be exposed to unforeseen problems; but then it is important to understand what could possibly generate these unforeseen problems. If we think of challenges that intelligent (animal) agents face in the real world (especially unforeseen ones), they very often come down to the fact that other agents are present (animals are in competition for resources, can be preys or predators, must reproduce, etc.).
@@
Based on such ideas, we propose to explore the space of Xent Games using a competitive vision which is close in spirit to the Novelty Search with Local Competition (NSLC) algorithms proposed in [LeSt11a, LeSt11b]. We can imagine a world where a number $N$ of agents play a (repeated) meta-game where each move is the generation of Xent Games. The meta-game design elements consist of roughly the following:
@@
  • Agents must at the same time play against the $k$ randomly sampled agents (again, in zero-sum mode) the games picked by the latter (on which they will have specialized, by symmetry).
If we think of a competitive landscape, agents will hence be incentivized to pick specialization in games that give them an edge over other agents’ games (because of the zero-sum nature of the games: these are resource allocation games), while at the same time picking games at which other agents’ skills are under-developed. This naturally suggests a competitive dynamics will take place, which should see the games evolve over time.
@@
This view of an evolving landscape generated by agents by a local game (or meta-game here) is for instance studied in [Lal2015]. The idea of using evolution-related ideas for artificial general intelligence has seen some uptick recently, see in particular [Ada23, Hug24, ZHLLC25]. Note that these ideas are also similar to the ideas of artificial curiosity and self-improvement [Schmi07, Schmi10, Clu19].
Leveraging such a dynamics as a ‘browsing mechanism’ (to use the analogy of Section 6.1.1) to build a game archive (the set of games visited) as an exploration dynamics to measure general abilities is appealing (as far as building a scope is concerned), as it provides a general idea of where ‘unforeseen new games’ could plausibly come emanate from; we could therefore use the list of visited games as a basis for a benchmark.
@@
While it can definitely be interesting to specify the details of the above meta-game, in the current paper we only use it as a means to inform the selection of relevant games to build appropriate Xent Measures. As such, it is in particular reasonable (assuming universality, see Section 6.2.3 below) to only assume that agents all use the same fine-tuned variants of the same base model $\mathcal{M}_{\mathcal{B}}$.
@@ 6.2.2. Key Features and Key Challenges of the Meta-Game
  • This suggests a view about general abilities formed by the aggregate of metastable equilibria: driven by the individual ‘initiatives’ that are selected on the basis being relevant between covering existing abilities and uncovering new directions. Hence the only way from the dynamics to not extend to some more broader (i.e. ‘uncaptured’) region of the game space is if there is no path of ‘meaningful’ games to it.
In spite of the appealing features listed above, a number of complex features are associated with the meta-game described in Section 6.2.1, make it in fact very under-specified and challenging to implement in practice:
  • Many unspecified details about the game are in fact difficult to set: how many players are involved throughout the steps, how are they different from one another, how much can they capitalize upon their previous choices, how are they making their choices, what information do they have about each other, etc?
  • The dynamics suggest an ‘intersubjective’ formulation of general capabilities: general capabilities move in some sense towards a direction resulting from what agents expect general abilities to look like. It can lead to a kind of very unstable set of games that ‘each believes the other believes, but no one truly believes it’.
@@
It should be obvious that determining a precise definitive answer to these questions, if one tries to model a real-world-inspired dynamics, is impossible. Ultimately, a way out is possible if the result of the dynamics, as far as measuring general capabilities is concerned, does not depend on these details: this is the view proposed in Section 6.2.3.
@@ 6.2.3. Universality: Do Exploration Details Matter?
Heuristically, there is a very simple reason why specific details should not matter: the whole point of any good measure of general abilities is that it should actually be fairly resilient to changes of the rules of games we use: agents with strong general capabilities will keep their strong performances on games that are known to be connected, while over-specialized agents are unlikely to get systematically lucky over a large, diverse enough family of games. Going further, any strong agent playing the meta-game should be playing in an environment where it can withstand any specific, computable choice of games: any such choice of games is in itself a possible play strategy, and any meta-game agent should choose games so that it can be resilient against such a specific strategy. Similarly, as suggested in Section 4.3.1), we expect the dependence of the measures of general abilities to be fairly independent of the precise setting (the $\mathcal{M},\mathcal{F},\mathcal{T}$ setting) associated with the transfer value functions (as long as they are ‘good’).
It is in fact reasonable to postulate the only relevant features are that the exploration covers a (1) connected and (2) diverse enough web of games, sampled in a (3) fair way. While the connectedness is actually easy to guarantee (by construction), it is important to make sure that the scope does not miss crucial games, i.e. games that are important to learn other games. As we will see in Section 6.3 below, a criterion that essentially implies this is the following:
@@
There exists a number $\alpha>0$ and a number $K\geq 1$ such that the space of Xent Games is $\alpha$-connected and such that for any game $G$ the maximal number $n$ of ‘escaping directions’, i.e. of games $G_{1},\ldots,G_{n}$ such that
@@
The above assumption suggests it is impossible to find very long sequences of ‘independent’ games that all capture some new meaningful information about any game: in other words, the number of different skills that have a substantial contribution to a game is finite. This natural assumption is for instance easily verified in the toy model introduced in Section 6.1.2 for $\alpha$ small enough and $K=\mathcal{O}\left(\left(1-\alpha\right)/\alpha\right)$.
From the Assumption 62, if there is an ‘undiscovered’ important game, there is at least a chain leading to it, and we will eventually find it, as it is impossible to ‘get lost’ in too many games starting from a game. Note that in practice, it is important to ensure that $K$ is reasonably small. From the Section 6.3, we propose a simple greedy algorithm aimed at extracting a sequence of games which arguably captures the same abilities as those discovered by a dynamics like the one suggested in Section 6.2.1.
@@ 6.3.2. Algorithm Specification
          If $\gamma\leq 1-\alpha_{2}$, add $G_{*}$ to $\Omega_{k}$.
@@ 6.3.3. Meta-Sampling Process
The objective of $\Lambda$ is quite clear: to provide the largest possible sample of games that allows the loop of the scope growth algorithm to run for as long as possible, in a way that maximally exhausts games. Given the structure of the Xent Games, the most useful practical implementations of such an algorithm involve sampling random codes in the XGL language from a dedicated model, trained to maximize the sampling quality. As discussed in Section 2.4.1, the Xent Game structure and the XGL is designed from the beginning with this goal in mind.
@@ 7.1. Summary
  • We investigate the question of what it means for an LLM to play Xent Games, propose evaluation procedures and introduce notions of playability and of transfer value.
@@
  • We propose to address the challenge of unbounded scope associated with measuring general capabilities (as opposed to limited to a finite scope) by relying on evolutionary dynamics principles, leading us to considerations close in philosophy to Quality-Diversity ideas.
@@ 7.2. Discussion
A number of important points in the approach above deserve some discussion.
@@ 7.2.1. Role of Judge Model
Throughout the paper, the Judge Model $\mathcal{J}$ is taken as a given. A central point is that for a sufficiently strong judge model, the space of implicit knowledge questions becomes large enough so that games can be played ‘ad infinitum’ (similarly to the fact that once one knows the rules of chess, one can keep improving at the game for a very long time). Still, improving the abilities of $\mathcal{J}$ or (perhaps counter-intuitively) allowing $\mathcal{J}$ to be ‘controllably weak’ (e.g. with certain prefix tokens) can lead to more sophisticated games; ultimately, the role of $\mathcal{J}$ is to provide an environment where abilities ought to be developed; as such, having high-quality data from the ‘external world’ contained in the training of the $\mathcal{J}$ model is fundamental, as the $\mathcal{J}$ model is the only connection between the external world and the model playing games.
@@ 7.2.2. Units of Measure
An important theme of Sections 4, 5, and 6 is to provide a clean interpretation of the quantities computed, in particular good units of measure. While the raw Xent Games values are measured in bits (which carries some useful problem-specific information), the measures shift towards relative units of work for the score normalizations with respect to a base model (comparing number of attempts) and the transfer value (such values are normalized by comparing token processing counts).
Ultimately this leads to a simple perspective on general capabilities which is in terms of units of work spared compared to specific replaying and training, assuming a common ‘reality’ defined by a model $\mathcal{J}$. As such, this provides a view of general capabilities that is continuous and does not contain any explicit thresholds for capabilities.
@@ 7.2.3. Connections
Ultimately, our construction relies upon a combination of ideas which appear in information theory, compositional game theory, transfer learning, search, benchmarking, and open-ended evolution. While some combinations of the above ideas are definitely well-studied, it is interesting to note that to the best of our knowledge, the connection between compositional game theory and transfer learning has not been studied at all; it is also interesting that the connection between open-ended evolution and general intelligence has seen a recent surge in interest [Ada23, XDCGZ24, ZHLLC25], but that no connection with the question of measuring intelligence has been provided yet.
@@ 7.3.1. Implementation
The most enticing practical challenge associated with the vision proposed in this paper is its implementation at scale: the determination and study of families of Xent Measures generated according to the ideas proposed in Section 6, including in particular efficient means to perform the game meta-sampling. In parallel, an interesting question is to investigate the potential of hand-crafted Xent Measures (i.e. for a finite set of hand-crafted games) as restricted benchmarks: in spite of their inherently narrow scopes, these already satisfy a number of interesting properties (in particular, not relying on a private dataset, and being generally hard to overfit on) that make them promising contenders to measure the abilities of models.

### v2

@@ 6.1.1. Web Exploration Intuition
If we assume that the space of Xent Games is $\alpha$-connected for some $\alpha>0$, we can think of it as an (infinite) graph and we can put a downstream link from $G_{1}$ to $G_{2}$ if $G_{1}\rightmoon_{\alpha}G_{2}$ and an upstream link if $G_{1}\leftmoon_{\alpha}G_{2}$. We can hence leverage this structure to think of the space of games as a kind of ‘inter-game web’ and use this vision as a guiding principle to explore it like a (very) complex network.
Note that unlike the internet case (where listing all the outgoing links from a page is trivial), it is a priori very non-trivial to list all the upstream and downstream links from a given game; however, we can think that in practice, if we have a finite list of games $\Sigma$, it is definitely possible to determine which ones are $\alpha$-linked and which ones are not, thus equipping $\Sigma$ with that structure.
@@
Following this principle, it is tempting to think of the exploration of the space of games as being similar to ‘random web browsing’ (following the philosophy of the PageRank algorithm [PBRW99]). If in some sense, one wants to find a ‘good sample’ of the internet, one can browse randomly and save the visited pages on the fly. Depending on the application, we may even move for long enough so that our visit becomes independent from the starting point (note also that the longer we ‘walk’, the more biased towards the more connected nodes we get). From this, we could, in principle, use various heuristics (e.g. the PageRank algorithm) to rank the games that we care about for a target application.
@@ 6.1.2. Toy Model
  • For any pair of games $G,\tilde{G}$, corresponding to $\left(g_{1},g_{2},\ldots,g_{n},\ldots\right)$ and $\left(\tilde{g}_{1},\tilde{g}_{2},\ldots,\tilde{g}_{n},\ldots\right)$, we define $\mathcal{V}_{G}\left(\tilde{G}\right)$ as $\frac{1}{\|\tilde{G}\|_{1}}\sum_{i=1}^{\infty}\min\left(\tilde{g}_{i},g_{i}\right)$, i.e. which fraction of $\tilde{g}$ skills are ‘captured’ by playing $g$ and, symmetrically, we define $\mathcal{V}_{\tilde{G}}\left(G\right)$ as $\frac{1}{\|G\|_{1}}\sum_{i=1}^{\infty}\min\left(g_{i},\tilde{g}_{i}\right)$.
@@
This simply suggests the (naive, but useful) view that games are associated with skills that are entries in a vector, and that learning these games amounts to acquiring skills proportionally to the vector of skills.
@@ 6.2.1. Meta-Game: Informal Evolution Dynamics Idea
In a sense, the question of measuring ‘general capabilities’ of models comes down to a motivation: why do we want to evaluate general abilities in the first place? A plausible answer is because agents will be exposed to unforeseen problems; but then it is important to understand what could possibly generate these unforeseen problems. If we think of challenges that intelligent (animal) agents face in the real world (especially unforeseen ones), they very often come down to the fact that other agents are present (animals are in competition for resources, can be preys or predators, must reproduce, etc.).
@@
Based on such ideas, we propose to explore the space of Xent Games using a competitive vision which is close in spirit to the Novelty Search with Local Competition (NSLC) algorithms proposed in [LeSt11a, LeSt11b]. We can imagine a world where a number $N$ agents play a (repeated) meta-game where each move is the generation of Xent Games. The meta-game design elements consist of roughly the following:
@@
  • Agents must at the same time play against the $k$ randomly sampled agents (again, in zero-sum mode) with the games picked by the latter (on which they will have specialized, by symmetry).
If we think of the competitive landscape, agents will hence be incentivized to pick specialization in games that give them an edge over other agents’ games (because of the zero-sum nature of the games: these are resource allocation games), while at the same time picking games at which other agents’ skills are under-developed. This naturally suggests a competitive dynamic will take place, which should see the games evolve over time.
@@
This view of an evolving landscape generated by agents in a local game (or meta-game here) is for instance studied in [Lal2015]. The idea of using evolution-related ideas for artificial general intelligence has seen some uptick recently, see in particular [Ada23, Hug24, ZHLLC25]. Note that these ideas are also similar to the ideas of artificial curiosity and self-improvement [Schmi07, Schmi10, Clu19].
Leveraging such a dynamic as a ‘browsing mechanism’ (to use the analogy of Section 6.1.1) to build a game archive (the set of games visited) used to measure general abilities is appealing (as far as building a scope is concerned), as it provides a general idea of where ‘unforeseen new games’ could plausibly come from; we could therefore use the list of visited games as a basis for a benchmark.
@@
While it can definitely be interesting to specify the details of the above meta-game, in the current paper we only use it as a means to inform the selection of relevant games to build appropriate Xent Measures. As such, it reasonable (assuming universality, see Section 6.2.3 below) to only assume that agents all use the same fine-tuned variants of the base model $\mathcal{M}_{\mathcal{B}}$.
@@ 6.2.2. Key Features and Key Challenges of the Meta-Game
  • This suggests a view about general abilities formed by the aggregate of metastable equilibria: driven by the individual ‘initiatives’ that are selected on the basis of being relevant between covering existing abilities and uncovering new directions. Hence, from the meta-game dynamics, the only for agents to not extend to broader (i.e. ‘uncaptured’) regions of the game space is if there are no paths of ‘meaningful’ games to them.
In spite of the appealing features listed above, a number of complex features associated with the meta-game described in Section 6.2.1 make it very under-specified and challenging to implement in practice:
  • Many unspecified details about the game are, in fact, difficult to set: how many players are involved throughout the steps, how are they different from one another, how much can they capitalize upon their previous choices, how are they making their choices, what information do they have about each other, etc?
  • The dynamics suggest an ‘intersubjective’ formulation of general capabilities: general capabilities move in some sense towards a direction resulting from what agents expect general abilities to look like. It can lead to a very unstable set of games that ‘each believes the other believes, but no one truly believes it’.
@@
It should be obvious that determining a precise, definitive answer to these questions, if one tries to model real-world-inspired dynamics, is impossible. Ultimately, a way out is possible if the result of the dynamics, as far as measuring general capabilities is concerned, does not depend on these details: this is the view proposed in Section 6.2.3.
@@ 6.2.3. Universality: Do Exploration Details Matter?
Heuristically, there is a very simple reason why specific details should not matter: the whole point of any good measure of general abilities is that it should actually be fairly resilient to changes of the rules of the games we use. Agents with strong general capabilities will keep their strong performances on games that are known to be connected, while over-specialized agents are unlikely to get systematically lucky over a large, diverse enough family of games. Going further, any strong agent playing the meta-game should be playing in an environment where it can withstand any specific, computable choice of games: any such choice of games is in itself a possible play strategy, and any meta-game agent should choose games so that it can be resilient against such a specific strategy. Similarly, as suggested in Section 4.3.1, we expect the dependence of the measures of general abilities to be fairly independent of the precise setting (the $\mathcal{M},\mathcal{F},\mathcal{T}$ setting) associated with the transfer value functions (as long as they are ‘good’).
It is, in fact, reasonable to postulate the only relevant features are that the exploration covers a (1) connected and (2) diverse enough web of games, sampled in a (3) fair way. While the connectedness is actually easy to guarantee (by construction), it is important to make sure that the scope does not miss crucial games, i.e. games that are important to learn other games. As we will see in Section 6.3 below, a criterion that essentially implies this is the following:
@@
There exists a number $\alpha>0$ and a number $K\geq 1$ such that the space of Xent Games is $\alpha$-connected and such that for any game $G$, the maximal number $n$ of ‘escaping directions’, i.e. of games $G_{1},\ldots,G_{n}$ such that
@@
The above assumption suggests it is impossible to find very long sequences of ‘independent’ games that all capture some new meaningful information about any game: in other words, the number of different skills that have a substantial contribution to a game is finite. This natural assumption is, for instance, easily verified in the toy model introduced in Section 6.1.2 for $\alpha$ small enough and $K=\mathcal{O}\left(\left(1-\alpha\right)/\alpha\right)$.
From the Assumption 62, if there is an ‘undiscovered’ important game, there is at least one chain leading to it and we will eventually find it, as it is impossible to ‘get lost’ in too many games starting from a game. Note that in practice, it is important to ensure that $K$ is reasonably small. From the Section 6.3, we propose a simple greedy algorithm aimed at extracting a sequence of games which arguably captures the same abilities as those discovered by dynamics like the one suggested in Section 6.2.1.
@@ 6.3.2. Algorithm Specification
          If $\gamma\leq 1-\alpha_{2}$: add $G_{*}$ to $\Omega_{k}$.
@@ 6.3.3. Meta-Sampling Process
The objective of $\Lambda$ is quite clear: to provide the largest possible sample of games that allows the loop of the scope growth algorithm to run for as long as possible, in a way that maximally exhausts games. Given the structure of the Xent Games, the most useful practical implementations of such an algorithm involve sampling random codes in the XGL language from a dedicated model, trained to maximize the sampling quality. As discussed in Section 2.4.1, the Xent Game structure and the XGL are designed from the beginning with this goal in mind.
@@ 7.1. Summary
  • We investigate the question of what it means for an LLM to play Xent Games, propose evaluation procedures, and introduce notions of playability and of transfer value.
@@
  • We propose to address the challenge of unbounded scope associated with measuring general capabilities (as opposed to limiting to a finite scope) by relying on evolutionary dynamic principles, leading us to considerations close in philosophy to Quality-Diversity ideas.
@@ 7.2. Discussion
A number of important points in the approach above deserve some discussion:
@@ 7.2.1. Role of Judge Model
Throughout the paper, the judge model $\mathcal{J}$ is taken as a given. A central point is that for a sufficiently strong judge model, the space of implicit knowledge questions becomes large enough so that games can be played ‘ad infinitum’ (this is similar to the fact that once one knows the rules of chess, one can keep improving at the game for a very long time). Still, improving the abilities of $\mathcal{J}$ or (perhaps counter-intuitively) allowing $\mathcal{J}$ to be ‘controllably weak’ (e.g. with certain prefix tokens) can lead to more sophisticated games. Ultimately, the role of $\mathcal{J}$ is to provide an environment where abilities ought to be developed. As such, having high-quality data from the ‘external world’ contained in the training of the $\mathcal{J}$ model is fundamental, as the $\mathcal{J}$ model is the only connection between the external world and an agent playing games.
@@ 7.2.2. Units of Measure
An important theme of Sections 4, 5, and 6 is to provide a clean interpretation of the quantities computed, in particular specifying good units of measure. While the raw Xent Games values are measured in bits (which carries some useful problem-specific information), the measures shift towards relative units of work for the score normalizations with respect to a base model (comparing number of attempts) and the transfer value (such values are normalized by comparing token processing counts).
Ultimately, this leads to a simple perspective on general capabilities which is specified in terms of units of work spared compared to specific replaying and training, assuming a common ‘reality’ defined by a model $\mathcal{J}$. As such, this provides a view of general capabilities that is continuous and does not contain any explicit thresholds for capabilities.
@@ 7.2.3. Connections
Ultimately, our construction relies upon a combination of ideas which appear in information theory, compositional game theory, transfer learning, search, benchmarking, and open-ended evolution. While some combinations of the above ideas are definitely well-studied, it is interesting to note that, to the best of our knowledge, the connection between compositional game theory and transfer learning has not been studied at all; it is also interesting that the connection between open-ended evolution and general intelligence has seen a recent surge in interest [Ada23, XDCGZ24, ZHLLC25], but that no connection with the question of measuring intelligence has been provided yet.
@@ 7.3.1. Implementation
The most enticing practical challenge associated with the vision proposed in this paper is its implementation at scale: this involves in particular the determination and study of families of Xent Measures generated according to the ideas proposed in Section 6, including, in particular, efficient means to perform the game meta-sampling and hyperparameter selection (e.g. the $\alpha$ values).
In parallel, an interesting question is to investigate the potential of hand-crafted Xent Measures (i.e. for a finite set of hand-crafted games) as restricted benchmarks: in spite of their inherently narrow scopes, these already satisfy a number of interesting properties (in particular, of not relying on a private dataset, and being generally hard to overfit on) that make them promising contenders to measure the abilities of models in a trusted and credible fashion.
7.3.2. Connection with Human-Centric Tasks
While the questions raised in Section 1.2 to motivate the investigation of the implicit knowledge have led us to games that are often directly inspired by an attempt to answer these questions (as illustrated in Section 2.3), to make the solutions to the games directly relevant to human-centric questions (e.g. to make good solutions of reverse prompting game 14 useful as human-directed summaries) likely requires some engineering (e.g. some appropriate pre-prompting or fine-tuning of the judge model focus on certain modes of expression, adding constraints, etc.).
From a capability measure perspective, by the very design of Xent game space (and the transfer learning considerations), it is reasonable to expect that models that display competence at Xent games will be capable to find solutions to their human-centric counterparts; at the same time, formulating precisely what the latter are is an important question.